[Brahms-l] Shift report 20050315 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue Mar 15 2005 - 15:56:28 EST
Shifter leader was JHL
Shifters were 


Beam:1.5 hour beam for physics (8:00- 9:30 am).
The beam for the rest of the shift was for 
machine developments at the injection energy.

Data: FS@8deg 861B program done: 400K trigger6
      MRS@45deg 700A: so far ~2M trigger3 
      Some data taken at injection energy: 13800-13804

Plan: FS@8deg 1219A: Till tomorrow's access

No detector/DAQ problems

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Mar 15 16:00:11 2005

This archive was generated by hypermail 2.1.8 : Tue Mar 15 2005 - 16:00:23 EST